1. Find a linear differential operator that annihilates the given function. (Use D for the differential operator.)

e−x + 5xex − x2ex

2. Find a linear differential operator that annihilates the given function. (Use D for the differential operator.)

cos 6x
1
Expert's answer
2020-06-29T18:43:08-0400

 1)A linear differential operator g(D) is called annihilates the function f(x) iff "g(f(x)) = 0" where "D = \frac{d}{dx}" .

Given "f(x)=e^{\u2212x} + 5xe^x \u2212 x^2e^x" _____________(1)

We known "(D-A)^n" annihilates"C_1e^{Ax}+C_2xe^{Ax}+C_3x^2e^{Ax}+ \u2026+ C_{n-1}x^{n-1}e^{Ax}" .

Hence, "(D+1)" annihilates "e^{-x}" and "(D-1)^3" annihilates "5xe^x-x^2e^x" .

Hence, "(D+1)(D-1)^3" annihilates "f(x)" .


2) Given"f(x) = cos(6x)" .

"\implies Df = -6sin(6x) \implies D^2f = -36 cos(6x) \\
\implies (D^2+36)f(x) = 0"

Thus, "(D^2+36)" annihilates "cos(6x)" .